OpenStudy (anonymous):

i need help on this question....34 * -4/5 = ??? Show work please help im really tired :(

OpenStudy (anonymous):

34 * (-4/5) = (-4)(34) / 5 = -136/5

OpenStudy (anonymous):

Wow your SUPER, SUPER SMART!!!

OpenStudy (anonymous):

-40/77 * (-44)=???

OpenStudy (anonymous):

= (-44x-40) / 77 = 1760 / 77 = 160/7

OpenStudy (anonymous):

Last question....

OpenStudy (anonymous):

You want to cut a bookmark that is 6 inches long and 2 3/8 inches wide from a 13 inch piece of paper.How much paper do i have left?